OpenStudy (anonymous):

A 1000 Ω resistor is in a circuit with 30.0 mA flowing through it. A voltmeter with internal resistance 1000 Ω is used to measure the voltage across the resistor. What is the reading from the voltmeter if 30.0 mA still flows in the circuit? 30.0 V 15.0 V 3.00 V 27.0 V 29.9 V

OpenStudy (anonymous):

The answer is 15.0V The explanation is as follows. Since the voltmeter is attached across the resistor, the arrangement is parallel. In a Parallel arrangement, The combined resistance is now 500ohms. Again, using V=IR V=(0.03)(500) V=15V--------------------------*Answer In a parallel arrangement, the voltage supplied remains constant to all units.

OpenStudy (anonymous):

are u sure ?

OpenStudy (anonymous):

I'm pretty positive. Would anyone mind backing up my reply?

OpenStudy (anonymous):

Attempt it yourself and compare answers. Use the formula V=IR, and assess the information as the only thing being changed here is the resistance. Apply that change to find out the effect it has on the voltage.

OpenStudy (anonymous):

pls @Areesha. ID how do you get the combined resistance to be 500ohms

OpenStudy (anonymous):

Alright, firstly voltmeters are always attached in a parallel arrangement so the combined rsistance of both would not simply be added. The formula for total resistance of resistors attached in a parallel arrangement is as follows: 1/RT= 1/R1 + 1/R2 + 1/R3 etc..